What is the ΔH° for the following reaction (in kJ/mol of reaction exactly as written)?

2 Al (s) + Fe2O3 (s) → Al2O3 (s) + 2 Fe (s)

To solve, use Appendix II in your textbook (containing standard heats of formation).
1
Expert's answer
2019-10-18T08:06:58-0400

Standard enthalies of formation of "Fe_2O_3" = "-822 \ kJ \ mol^{-1}"

Standard enthalies of formation of "Al_2O_3=-1669\ kJ\ mol^{-1}"

Now ,


"\Delta H^{o}_r=\Sigma(\Delta H^o_{products}-\Delta H^o_{reactants})""\Delta H^o_r=-1669-(-822)=-847\ kJ\ mol^{-1}"